### v2

# Semantics Prompting Data-Free Quantization for Low-Bit Vision Transformers

Abstract
Data-free quantization (DFQ), which facilitates model quantization without real data to address increasing concerns about data security, has garnered significant attention within the model compression community. Recently, the unique architecture of vision transformers (ViTs) has driven the development of specialized DFQ techniques. However, we observe that the synthetic images from existing methods suffer from the deficient semantics issue compared to real images, thereby compromising performance. Motivated by this, we propose SPDFQ, a Semantics Prompting Data-Free Quantization method for ViTs. First, SPDFQ incorporates Attention Priors Alignment (APA), which uses randomly generated attention priors to enhance the semantics of synthetic images. Second, SPDFQ introduces Multi-Semantic Reinforcement (MSR), which utilizes localized patch optimization to prompt efficient parameterization and diverse semantics in synthetic images. Finally, SPDFQ employs Softlabel Learning (SL), where soft learning targets are adapted to encourage more complex semantics and accommodate images augmented by MSR. Experimental results demonstrate that SPDFQ significantly outperforms existing methods. For instance, SPDFQ achieves a 15.52% increase in top-1 accuracy on ImageNet for W4A4 ViT-B [^1] The code will be made publicly available upon acceptance..
Figure 1: Feature visualization using $t$-SNE [63]. Each shape (circle, star, triangle) represents a category. The red dashed circles highlight the features of our APA and real images. The visualized features are activations extracted before the classification head of DeiT-S. The features from PSAQ-ViT [43] deviate significantly from those of the real images, suggesting deficient semantics. While the features from our APA are more closely aligned with those of real images, suggesting improved semantics.
1 Introduction
Vision transformers (ViTs) [19] have attracted widespread interest from both academia and industry [33, 26] due to their superior performance across various vision tasks [62, 7, 1, 52]. However, the substantial computational demands and high memory requirements of ViTs pose significant challenges for deployment in resource-limited environments [61, 27, 29, 75, 42, 51, 11]. To address this challenge, model quantization [36] has emerged as a promising solution, reducing model complexity by enabling a low-bit representation of weights and activations. Nevertheless, most existing quantization methods rely on access to the original training data, raising concerns regarding data privacy and security. This data dependency restricts their application in data-sensitive environments [6, 77, 24, 70].
To address this limitation, data-free quantization (DFQ) has been developed to quantize models without accessing real data, attracting considerable attention [69, 74, 14]. Most existing DFQ methods are tailored to convolutional neural networks (CNNs) and are not directly applicable to ViTs. These methods typically utilize batch normalization statistics (BNS), which capture the distribution of real data, to synthesize in-distribution synthetic data [6, 69, 74, 77]. However, BNS is unavailable for ViTs that use layer normalization (LN) to dynamically compute distribution statistics during inference [43]. Recently, several works have been proposed to accommodate the unique structure of ViTs [43, 44, 58, 16, 30]. For example, PSAQ-ViT [43] introduces patch similarity entropy (PSE) loss to optimize Gaussian noise towards usable synthetic images.
Nevertheless, we observe that the current synthetic images suffer from the deficient semantics issue. In Fig. 1, we visualize the penultimate features of the synthetic images using t-SNE [63]. It can be observed that the features of synthetic images generated by PSAQ-ViT diverge significantly from those of real images, suggesting that these synthetic images are semantically deficient. This observation is further supported by the quantitative results in Tab. 1, where the cosine similarity between synthetic images from PSAQ-ViT and real images is notably low. For high-bit quantization, where model capacity is largely retained [39], the performance degradation remains relatively minor even with deficient semantics images [43, 58]. However, in low-bit quantization, where model capacity is severely damaged and informative images are essential for recovering performance [77, 13], fine-tuning on these deficient semantics images leads to poor generalization to real datasets, resulting in limited performance gains.
Motivated by this finding, we propose a Semantics Prompting Data-Free Quantization method for ViTs, termed SPDFQ. The overall framework is depicted in Fig. 2. First, SPDFQ introduces Attention Priors Alignment (APA), which employs randomly generated priors to encourage high responses in attention maps, thereby explicitly promoting semantics in synthetic images. These priors are generated using the Gaussian Mixture Model (GMM) to simulate diverse attention patterns. APA effectively prompts semantics in synthetic images, as validated by both visualization and quantitative analyses. As shown in Fig. 1, features of APA exhibit a closer alignment to those of real images, indicating high semantics. The quantitative results in Tab. 1 further confirm that the semantics of APA are more similar to the real images. Additionally, SPDFQ incorporates Multi-Semantic Reinforcement (MSR) and Softlabel Learning (SL). MSR promotes efficient parameterization and diverse semantics in synthetic images by leveraging localized patch-wise optimization, while SL prompts more complex semantics in synthetic images and accommodates images augmented by MSR through the use of soft learning targets.
Experimental results across various ViT models demonstrate that SPDFQ presents substantial performance improvements. For example, SPDFQ achieves a 15.52% increase in top-1 accuracy on the ImageNet dataset for the W4A4 ViT-B model.
2 Related Works
2.1 Vision Transformers
The great success of transformers in the natural language processing field has driven widespread attempts in the computer vision community to apply them to vision tasks [12, 65, 25]. ViT [19] is the pioneer that builds a transformer-based model to handle images, boosting the performance on the image classification task. DeiT [62] introduces an efficient teacher-student training strategy where a distillation token is employed to distill knowledge from the teacher model to the student model. Swin Transformers [50] builds an efficient and effective hierarchical model by introducing a shifted window-based self-attention mechanism. Other than the image classification task, the applications of ViTs also have broadened considerably, manifesting groundbreaking performance in object detection [7], image segmentation [9, 76], low-level vision [46], video recognition [55, 1], and medical image processing [60], etc. Nevertheless, the impressive performance of ViTs relies on a high number of parameters and significant computational overhead, preventing deployment in resource-constrained environments. Several recent efforts design lightweight ViTs, such as MobileViT [53], MiniVit [73], and TinyViT [66]. However, the model complexity is still unsatisfactory [43].
2.2 Network Quantization
Data-Driven Quantization. Model quantization reduces the complexity of neural networks by replacing full-precision weight and activation with the low-bit format. Data-driven quantization can be roughly divided into two categories: quantization-aware training (QAT) and post-training quantization (PTQ). QAT is compute-heavy since it re-trains the quantized model with the full training data to retain performance [40, 42, 38, 22, 48, 20, 67]. PTQ perform quantization with a tiny dataset and a reduced time overhead, harvesting widespread attention [3, 39]. The specific architecture of ViTs, such as LayerNorm and the self-attention module, urges distinct PTQ methods compared to CNNs [49, 18, 47, 42, 21, 79]. For example, Liu et al[51] develop a ranking loss to maintain the relative order of the self-attention activation. Unfortunately, both QAT and PTQ involve the original training data, causing concerns about data privacy and security issues in data-sensitive scenarios.
Data-Free Quantization. DFQ quantizes models without accessing real data [24, 14, 41, 2, 13, 57, 37, 32, 15]. Most previous DFQ methods focus on CNN, where the BNS can be adopted as the regularization term [6, 74]. However, BNS is infeasible for ViTs built on the LN. Recently, few efforts have been explored to accommodate ViTs [43, 44, 58, 16, 30]. PSAQ-ViT [43] introduces the first DFQ method for ViTs. They discover that Gaussian noise yields homogeneous patches, while the real image yields heterogeneous patches. Thus, patch similarity entropy (PSE) loss is proposed to optimize the Gaussian noise towards real-like images by making them showcase heterogeneous patches. Based on PSAQ-ViT, PSAQ-ViT V2 [44] further introduces a adversarial learning strategy [23]. [58] incorporates contrastive learning and proposes an iterative generation-quantization PTQ-based DFQ method. [30] proposes a sparse generation method to remove noisy and hallucination backgrounds in synthetic images.
Figure 2: Framework of the proposed SPDFQ. Attention Priors Alignment (APA) employs randomly generated attention priors to explicitly promote semantics in synthetic images. Multi-Semantic Reinforcement (MSR) prompts the different parts of synthetic images with various semantics. Softlabel Learning (SL) leverages soft learning to prompt more complex semantics in synthetic images.
3 Method
3.1 Preliminaries
3.1.1 Quantizers
We employ the linear quantizer for all weights and activations, except for the attention scores, which use a log2 quantizer to handle highly non-negative and uneven values [45, 47, 21]. For the the linear quantizer, given a full-precision input $\mathbf{x}$ and bit-width $b$, the quantized value $\mathbf{x}_{q}$ and the de-quantized value $\bar{\mathbf{x}}$ are computed as follows:
$\mathbf{x}_{q}$ $=\text{clip}\left(\left\lfloor\frac{\mathbf{x}}{\Delta}\right\rceil+z,0,2^{b}-1\right),\bar{\mathbf{x}}=\Delta\cdot\left(\mathbf{x}_{q}-z\right),$ (1)
[…]
$\mathbf{x}_{q}$ $=\text{clip}\left(\left\lfloor-\log_{2}\frac{\mathbf{x}}{\Delta}\right\rceil,0,2^{b}-1\right),\bar{\mathbf{x}}=\Delta\cdot 2^{-\mathbf{x}_{q}}.$ (2)
3.1.2 Data Synthesis
DFQ methods parameterize synthetic images and optimize them toward real-like images with a pre-trained full-precision model $F$. Given a image $\tilde{\bm{I}}$ initialized from Gaussian noise, the one-hot loss [68] is introduced to prompt label-related semantics:
$\mathcal{L}^{\text{OH}}(\tilde{\bm{I}})=CE(F(\tilde{\bm{I}}),c)$ (3)
where $CE(\cdot,\cdot)$ represents the cross entropy, $c$ is a random class label, and $F(\cdot)$ returns the predicted probability for image $\tilde{\bm{I}}$.
Moreover, total variance (TV) [70] loss is a smoothing regularization term to improve the image quality:
[…]
where $\nabla\tilde{\bm{I}}(\tau_{1},\tau_{2})$ denotes the gradient at $\tilde{\bm{I}}$ at $(\tau_{1},\tau_{2})$.
To perform DFQ for ViTs, PSAQ-ViT [43] proposes patch similarity entropy (PSE) loss. It first compute patch similarity $\bm{\Gamma}_{l}[i,j]=\frac{u_{i}\cdot u_{j}}{||u_{i}||||u_{j}||}$, where $u_{i},u_{j}$ are feature vectors of MHSA outputs in $l$-th block and |||||$\cdot$||| | ⋅ | | denotes the $l_{2}$ norm. Then, it estimates the density function $\hat{f_{l}}(x)=\frac{1}{Mh}\sum_{m=1}^{M}K\left(\frac{x-x_{m}}{h}\right)$, where $K(\cdot)$ is a normal kernel, $h$ is the bandwidth, and $x_{m}$ is the kernel center derived from $\bm{\Gamma}_{l}$. Finally, the PSE loss is defined as:
[…]
where $L$ is the block number of the model.
3.2 Observations
Although existing DFQ methods of ViTs have made some progress, a significant performance gap remains between synthetic and real images. For example, as shown in Tab. 2, the W4A4 ViT-B fine-tuned on real images yields 68.16%, whereas PSAQ-ViT only achieves 36.32%. To dive into a deeper analysis, we visualize the features of the synthetic images before the classification head using t-SNE [63] in Fig. 1. Generally, the penultimate feature represents the semantics of the input [72, 4, 54, 71].
It is evident that features of PSAQ-ViT diverge significantly from those of real samples, indicating that the synthetic images fail to capture the real semantics distribution. We refer to this phenomenon as the deficient semantics issue. In Tab. 1, we quantitatively measure the semantics using the average cosine similarity (ranging from $-1$ to $1$). We also report the intra-class similarity within real images as an approximate upper bound for comparison. These results further highlight the deficient semantics of PSAQ-ViT. For example, for Class 1, the intra-class similarity within real images is 0.68, whereas PSAQ-ViT only presents 0.44, suggesting a low similarity between synthetic and real images. Similarly, for Class 3, while the real images exhibit a similarity of 0.41, PSAQ-ViT achieves only 0.31.
Both visualization and quantitative results reveal a pronounced deficient semantics. Consequently, quantized models fine-tuned on these semantically deficient synthetic images cannot well generalize to the real dataset, resulting in limited performance.
Classes Method Real PSAQ-ViT [43] APA (Ours)
Class 1 0.68 0.44 0.64
Class 2 0.32 0.26 0.31
Class 3 0.41 0.31 0.36
Table 1: Average cosine similarity of three randomly selected classes. For real images, the similarity is measured within the class itself, while for PSAQ-ViT and APA, the similarity is measured between synthetic and real images of the same class. The results show that APA achieves higher similarity than PSE loss, indicating improved semantics.
3.3 Solutions
In the following, we introduce the proposed SPDFQ to prompt semantics in synthetic images. As illustrated in Fig. 2, SPDFQ introduces three contributions: Attention Priors Alignment (APA), Multi-Semantic Reinforcement (MSR), and Softlabel Learning (SL).
3.3.1 Attention Priors Alignment
The first step of SFDFQ in prompting semantics leverages the self-attention mechanism of ViTs. Specifically, the self-attention mechanism extracts and integrates semantic information by allocating more attention to tokens rich in semantics and less to tokens with minimal semantics [19, 28, 10]. Thus, if a synthetic image is semantically rich, its attention maps should exhibit high responses. Based on this characteristic, we propose Attention Priors Alignment (APA), which is designed to promote the semantics of synthetic images by explicitly encouraging high responses in images’ attention maps.
For a synthetic image $\tilde{\bm{I}}$, we first extract its attention maps of the $h$-th head in the $l$-th block, denoted by $\mathbf{A}_{l,h}\in\mathbb{R}^{N\times N}$, by inputting the image into a pre-trained full-precision model, where $N$ represents the total number of tokens. In DeiT, the attention of the classification token toward other tokens serves as the indicator for semantic versus non-semantic parts [5]. Accordingly, we extract $\mathbf{A}^{c}_{l,h}\in\mathbb{R}^{1\times(N-1)}$ from $\mathbf{A}_{l,h}$, representing the attention of the classification token to all tokens except itself. We then randomly generate attention priors $\tilde{\mathbf{A}}_{l,h}$, whose generation is detailed in the next part, and align $\mathbf{A}^{c}_{l,h}$ with $\tilde{\mathbf{A}}_{l,h}$ by:
$\mathcal{L}_{l,h}(\tilde{\bm{I}})=\text{MSE}(\mathbf{A}^{c}_{l,h}-\tilde{\mathbf{A}}_{l,h}),$ (6)
where MSE represents the mean squared error. For Swin models that do not use a classification token, we substitute $\mathbf{A}^{c}_{l,h}$ in Eq. 6 with the average attention map of all tokens [10]. As noted in [17], ViTs initially focus on all regions to capture low-level information in shallow blocks and gradually shift their focus toward semantic regions in deeper blocks to extract high-level semantic information. Leveraging this property, we selectively apply $\mathcal{L}^{\text{APA}}_{l,h}$ to deeper blocks, progressively aligning attention towards semantically relevant areas. The total APA loss is computed as a depth-weighted sum of the individual Eq. 6 across these deeper blocks:
$\mathcal{L}^{\text{APA}}(\tilde{\bm{I}})=\sum_{l=S}^{L}\sum_{h=1}^{H}\frac{l}{L}\mathcal{L}_{l,h}(\tilde{\bm{I}}),$ (7)
where $S$ is a pre-given hyper-parameter denoting the start of deep blocks and are experimentally set $S=\frac{L}{2}$.
The proposed APA loss encourages synthetic images to contain semantically meaningful regions with high responses in attention maps, thereby prompting the semantics of the generated images.
Figure 3: Examples of generated attention priors.
Attention Priors Generation. To generate attention priors $\tilde{\mathbf{A}}_{l,h}$, we employ GMM for it is the most commonly used distribution with high flexibility. ViTs utilize different attention heads to capture diverse patterns and learn varied, informative representations [8]. Thus, we use distinct GMMs for each head. Note that the goal here is not to replicate real attention maps precisely, but to generate simulated attention priors. These priors are used to explicitly encourage semantically meaningful regions within synthetic images, prompting high responses in attention maps to enhance semantics.
In particular, we first initialize an all zero matrix $\tilde{\mathbf{P}}\in\mathbb{R}^{H\times W}$, where $H=W=\sqrt{N-1}$ for DeiT, $H=W=\sqrt{N}$ for Swin. For example, for DeiT-S, the $H=W=\sqrt{196}=14$. Then, we generate $k$ two-dimensional Gaussian distributions, where $k$ is randomly sampled from $1\sim K_{APA}$ and $K_{APA}$ is set to $5$ in all experiments. Each Gaussian has its mean and covariance [^2] The pseudo code is detailed in supplementary materials.. Consequently, the matrix element at the $i$-th row and $j$-th column, $\tilde{\mathbf{P}}[i,j]$, is determined by:
$\tilde{\mathbf{P}}[i,j]=\max_{m=1,\cdots,k}\mathbf{G}^{m}[i,j].$ (8)
Then, $\tilde{\mathbf{P}}$ is normalized by:
$\tilde{\mathbf{P}}_{n}=\frac{\tilde{\mathbf{P}}}{\sum\tilde{\mathbf{P}}}*(1-x).$ (9)
Here, for DeiT that incorporates the classification token, $x$ is randomly sampled from a uniform distribution $U(0,1)$, representing the proportion of the attention score that the classification token allocates to itself. For Swin, which does not use a classification token, $x$ is set to 0. Finally, $\tilde{\mathbf{P}}_{n}$ is flatten to match the dimensionality:
$\tilde{\mathbf{A}}_{l,h}=\text{flatten}(\tilde{\mathbf{P}}_{n}).$ (10)
Fig. 3 displays examples of generated attention priors. Although simple, APA effectively enhances semantics, as validated by both visualization and quantitative results. As shown in Fig. 1, compared to PSAQ-ViT, features after applying the APA loss are closer to the real images, indicating improved semantics. The quantitative results in Tab. 1 further support that APA achieves higher semantics. For instance, in Class 1, the intra-class similarity of PSAQ-ViT is 0.44, whereas APA achieves an improved similarity of 0.64.
3.3.2 Multi-Semantic Reinforcement
Current methods [43, 44, 58] optimize the entire synthetic image, restricting it to capture only the semantics of a single object type. However, this has two limitations. First, real images naturally exhibit low-rank properties, such as spatial similarity between nearby pixels [31]. Optimizing the entire images overlooks this structural regularity, easily leading to synthetic images with noisy or spurious regions that convey meaningless or misleading semantics [30]. This results in inefficient parameterization, where semantic contents available for model training are limited [34]. Second, real images often contain multiple objects with varied semantics. Constraining synthetic images to a single object type limits their semantic richness and diversity.
Motivated by the above analysis, we propose Multi-Semantic Reinforcement (MSR), which leverages localized patch optimization to promote efficient parameterization and diverse semantics. Specifically, for a synthetic image $\tilde{\bm{I}}$, instead of feeding only the entire image, we also feed its patches and optimize them individually. Initially, we select $m$ non-overlapping patches, where $m$ is chosen randomly from the set $\{1,2,\ldots,K_{MSR}\}$, with $K_{MSR}$ set to 4 in all experiments. These $m$ patches are then cropped and resized to match the model’s input dimensions:
$\{\tilde{\bm{I}}^{i}_{MSR}\}_{i=1,\cdots,m}=\text{resize}\big{(}\text{crop}_{m}(\tilde{\bm{I}})\big{)},$ (11)
where $\text{crop}_{m}(\cdot)$ crops $m$ non-overlapping patches from input, and $\text{resize}(\cdot)$ is the resize function. Each patch, denoted as $\tilde{\bm{I}}^{i}_{MSR}$, is treated as a new image with an assigned semantic target $c^{i}$. For simplicity, the loss function used to learn label-related semantics in both $\tilde{\bm{I}}^{i}_{MSR}$ and $\tilde{\bm{I}}$ is presented in Sec. 3.3.3. Note that the gradient is backpropagated only to update the corresponding patch in the original image, leaving the rest of the image unaffected.
By leveraging localized patch optimization, MSR effectively promotes efficient parameterization and diverse semantics in synthetic images. First, MSR reduces noisy and non-semantic regions and provides more distinct semantic samples, i.e., $\{\tilde{\bm{I}}^{i}_{MSR}\}_{i=1,\cdots,m}$, for model training, achieving efficient parameterization. Second, by assigning different semantic targets to individual cropped patches, MSR transforms synthetic images into composites of multiple semantic objects rather than a single uniform entity, thereby enhancing overall semantic diversity. Unlike traditional cropping used in data augmentation, which aims to improve classification robustness by training with cropped patches labeled with the original class, MSR aims to prompt efficient parameterization and semantic diversity within synthetic images, ultimately enabling accurate data-free quantization.
3.3.3 Softlabel Learning
The one-hot loss (Eq. 3) optimizes synthetic images until their confidence for the target class approaches 1. However, this leads to overfitting on a specific class, limiting semantic richness [77]. Moreover, Eq. 3 only learns the semantics of the target class. making it unsuitable for $\tilde{\bm{I}}$ under MSR, as its patches $\{\tilde{\bm{I}}^{i}_{MSR}\}_{i=1,\cdots,m}$ contain distinct semantics.
To address this, we propose Softlabel Learning (SL), which softens the learning objective to promote more complex semantics and accommodate images augmented by MSR. Specifically, we first sample $Z\in\mathbb{R}^{C}\sim U(0,1)$, then modify its values by:
$\begin{cases}Z[c^{i}]\sim U(\epsilon_{1},\epsilon_{2}),&\text{for }\tilde{\bm{I}}^{i}_{MSR},\\ Z[c^{1},\ldots,c^{m}]\sim U(\epsilon_{1},\epsilon_{2}),&\text{for }\tilde{\bm{I}},\end{cases}$
where $U(\epsilon_{1},\epsilon_{2})$ denotes the uniform distribution over the interval $[\epsilon_{1},\epsilon_{2}]$, $m$ is the number of patches determined in MSR, and $\epsilon_{1}$ and $\epsilon_{2}$ control the softness, both empirically set consistently to 5 and 10 in all experiments. The soft target is defined as $T_{s}=\text{softmax}(Z)$, and SL loss is:
${\cal L}^{\text{SL}}(\tilde{\bm{I}}/\tilde{\bm{I}}^{i}_{MSR})=SCE\big{(}F(\tilde{\bm{I}}/\tilde{\bm{I}}^{i}_{MSR}),T_{s}\big{)},$ (12)
where $SCE(\cdot,\cdot)$ is soft cross entropy and $F(\cdot)$ returns the predicted probability for its input.
For $\tilde{\bm{I}}^{i}_{MSR}$, SL aligns its predicted results with a “soft label” rather than a traditional one-hot label, mitigating overfitting and promoting the generation of semantically complex content. For $\tilde{\bm{I}}$, SL provides a suitable target for learning images with multiple semantic targets. Unlike the Soft Inception loss [77], which focuses only on the primary class probability, SL imposes a distribution across all classes, offering a comprehensive learning objective. Additionally, SL is compatible with MSR, where images contain multiple semantic targets. In contrast, the Soft Inception loss is designed for single-class scenarios and is therefore incompatible.
3.4 Overall Pipeline
The overall pipeline consists of two stages: data synthesis and quantized network learning. The first stage synthesizes synthetic images. The second stage fine-tunes the quantized model using the generated synthetic images.
3.4.1 Data Synthesis
In the data synthesis stage, we combine the proposed APA loss of Eq. 7, SL loss of Eq. 12, and TV loss of Eq. 4 to formulate the objective function as follows:
$\mathcal{L}_{G}(\tilde{\bm{I}})$ $=\alpha_{1}\mathcal{L}^{\text{APA}}(\tilde{\bm{I}})+{\cal L}^{\text{SL}}(\tilde{\bm{I}})+0.05\mathcal{L}^{\text{TV}}(\tilde{\bm{I}}).$ (13)
where $\alpha_{1}$ is hyperparameters and is determined by grid-search. Note that the weight of TV loss is fixed to 0.05, following [43], to avoid a cumbersome hyperparameter search.
3.4.2 Quantized Network Learning
Recently DFQ methods have introduced the PTQ methods in learning quantized models due to their advantages of speed, memory efficiency, and performance [32, 58]. Thus, following the success of [39, 64], we fine-tune the quantized network block-wisely. Specifically, denote $\mathbf{X}_{l}$ as the outputs of the $l$-th block of the full-precision model, and $\bar{\mathbf{X}}_{l}$ represent outputs of the quantized counterpart. The reconstruction loss is defined as:
$\mathcal{L}_{l}=\|\mathbf{X}_{l}-\bar{\mathbf{X}}_{l}\|_{2}.$ (14)
Here, $\mathcal{L}_{l}$ is only backward to update weights within $l$-th block. Note that for a fair comparison, all compared methods adopt the same quantized network learning stage.
4 Experiment
4.1 Implementation Details
Models and Datasets. We evaluate the performance of the proposed SPDFQ across various ViTs on the image classification task, including ViT-S/B [19], DeiT-T/S/B [62], and Swin-S/B [50]. The dataset is ImageNet [59], which contains 1.28M training images and 50K validation images across 1,000 classes. The pre-trained models are downloaded from the timm library.
Model W/A Real Gaussian noise PSAQ-ViT [43] PSAQ-ViT V2 [44] SMI [30] SPDFQ (Ours)
ViT-S (81.39) 4/4 66.57 6.02 47.24 41.53 24.3329.41 50.32
5/5 76.69 36.77 71.59 68.41 61.3365.19 74.31
6/6 79.46 61.20 77.20 74.76 72.9572.46 78.40
ViT-B (84.54) 4/4 68.16 0.15 36.32 26.32 35.2719.67 51.84
5/5 79.21 4.16 68.48 67.95 67.5357.13 70.70
6/6 81.89 55.18 76.65 71.87 76.3369.82 79.16
DeiT-T (72.21) 4/4 56.60 17.43 47.75 30.20 30.1413.18 52.06
5/5 67.09 43.49 64.10 55.16 56.4439.35 66.41
6/6 69.81 56.23 68.37 62.77 64.0344.39 69.73
DeiT-S (79.85) 4/4 68.46 20.89 58.28 45.53 42.7711.71 62.29
5/5 75.06 41.06 71.90 63.14 62.8829.13 74.06
6/6 77.87 65.63 75.85 68.85 71.6537.69 77.31
DeiT-B (81.85) 4/4 77.07 47.20 71.75 66.43 65.3359.04 72.17
5/5 79.86 65.46 78.45 76.77 76.7475.33 78.72
6/6 80.90 62.79 80.00 79.22 78.8177.66 80.15
Swin-S (83.20) 4/4 78.12 31.92 73.19 65.55 65.85 74.74
5/5 80.51 52.10 78.15 74.37 75.41 79.56
6/6 80.60 65.66 79.74 78.50 78.25 80.56
Swin-B (85.27) 4/4 78.80 30.14 71.84 67.42 65.23 76.42
5/5 82.51 35.28 78.50 77.20 75.25 80.82
6/6 82.64 67.37 82.00 81.41 80.30 83.03
Table 2: Quantization results on ImageNet dataset, with top-1 accuracy (%) reported. The performance of the full-precision model is listed below the model name. “W/A” denotes the bit-width of weights/activations. “Real” refers to using real images. For SMI [30], we provide the performance of using dense (normal-sized numbers) and sparse (smaller-sized numbers) synthetic images, respectively. Note that for Swin models, we do not provide the results for sparse synthetic images as the sparse generation method of SMI is infeasible.
Comparison methods. We compare our SPDFQ against Gaussian noise, real images, and previous methods including SMI [30] and PSAQ-ViT [43] and its subsequent version, PSAQ-ViT V2 [44]. For a fair comparison, we generate synthetic images using their methods and apply our quantized network learning strategy. We use the official code for SMI and PSAQ-ViT to reproduce their images, while PSAQ-ViT V2 is re-implemented by us, as no official code is available. [^3] In supplementary materials, we present more comparisons with more methods on W8/A8 and W4/A8 settings.
Experimental settings. All experiments use the PyTorch framework [56] on a single NVIDIA 3090 GPU. In the data synthesis stage, synthetic images were initialized with standard Gaussian noise, with 32 synthetic images generated in total. The Adam optimizer [35] with $\beta_{1}=0.5$, $\beta_{2}=0.9$ is employed. The learning rate is 0.25 for Swin and 0.2 for others, with a total of 1,000 iterations. For all models, $K_{APA}$, $K_{MSR}$, $\epsilon_{1}$, and $\epsilon_{2}$ are respectively set to 5 and 4, 5, and 10, with these values are searched using W4A4 DeiT-S. The value of $\alpha_{1}$ was determined by grid search for each model. Specifically, $\alpha_{1}$ is set to 1e5 for DeiT-T/S, 1e4 for DeiT-B, 100 for ViT-B and Swin-B, 10 for Swin-S, and 1 for ViT-S. Although further searching hyperparameters may improve the performance, we use current settings since these already yield superior results. In the quantized network learning stage, the Adam optimizer with $\beta_{1}=0.9,\beta_{2}=0.999$ is employed. The weight decay is set to 0, and the initialized learning rate is 4e-5, adjusted by the cosine decay strategy. The total iteration is set to 100. Consistent with [45], quantization parameters were determined by searching for percentile values that minimize quantization error before and after quantization. We follow the two-stage optimization strategy in [78] to optimize the quantized network. The channel-wise quantizer is adopted for weights and the layer-wise quantizer for activations. All matrix multiplications in ViTs are quantized [44, 43, 45].
4.2 Quantization Results
The quantization results are presented in Tab. 2. The proposed SPDFQ demonstrates consistent improvements across various quantization bit-width configurations, particularly with low bit-width settings. Specifically, for ViT-S, SPDFQ improves the performance by 3.08% in the W4/A4 setting, 2.72% in the W5/A5 setting, and 1.20% in the W6/A6 setting. For ViT-B, SPDFQ achieves performance gains of 15.52% in the W4/A4 setting, 2.22% in the W5/A5 setting, and 2.51% in the W6/A6 setting. Results on DeiT also demonstrate the effectiveness of the proposed SPDFQ. For DeiT-T, SPDFQ shows a marked improvement by increasing top-1 accuracy by 4.31% in the W4/A4 setting, 2.31% in the W5/A5 setting, and 1.36% in the W6/A6 setting. For DeiT-S, SPDFQ enhances top-1 accuracy by 4.01% in the W4/A4 setting, 2.16% in the W5/A5 setting, and 1.46% in the W6/A6 setting. As for DeiT-B, SPDFQ enhances top-1 accuracy by 0.42% in the W4/A4 setting, 0.27% in the W5/A5 setting, and 0.15% in the W6/A6 setting. The quantization results of Swin-S/B also affirm the superiority of our SPDFQ in enhancing model accuracy under different quantization configurations. In particular, for Swin-S, the proposed SPDFQ increases the accuracy by 1.55% for the W4/A4 setting, 1.41% for the W5/A5 setting, and 0.82% for the W6/A6 setting, respectively. When it comes to Swin-B, the proposed SPDFQ increases the accuracy by 4.58% for the W4/A4 setting, 2.32% for the W5/A5 setting, and 1.03% for the W6/A6 setting, respectively.
4.3 Ablation Study
All ablation studies are conducted on the W4A4 DeiT-S.
APA MSR SL Acc. (%)
Baseline 51.73
60.26
50.75
52.02
61.58
60.51
56.08
62.29
Table 3: Influence of the proposed APA, MSR, and SL on accuracy. The baseline adopts the one-hot loss.
Analysis of APA, MSR, and SL. We first analyze the effectiveness of the proposed APA (Sec. 3.3.1), MSR (Sec. 3.3.2), and SL (Sec. 3.3.3). Experimental results are present in Tab. 3. It can be seen that when the APA and SL are individually added to synthesize fake images, the accuracy increases compared with the baseline. Notably, APA significantly boosts the baseline from 51.73% to 60.26%, supporting the analysis in Sec. 3.3.1 that demonstrates the APA’s effectiveness in enhancing semantics. In contrast, applying MSR alone slightly reduces accuracy from 51.73% to 50.75%, indicating that one-hot loss is unsuitable for synthetic images augmented with MSR. However, when both MSR and SL are applied, performance rises to 56.08%, suggesting that SL is more compatible with MSR than the one-hot loss. Combining APA with either MSR or SL further increases performance. For example, applying both APA and SL presents an accuracy of 60.51%. When all of the three strategies are applied, the best performance of 62.29% can be obtained.
Analysis of priors distribution. Tab. 5 showcases the results of using other distributions to formulate the attention priors. The unevenly distributed GMM and Laplace present comparable performance of 62.29% and 62.16%, respectively. Moreover, GMM provides a similar performance to the real’s, indicating it performs well in imitating the patterns of real images.
Analysis of $\alpha_{1}$, $K_{APA}$, and $K_{MSR}$. The $\alpha_{1}$ from Eq. 13 balance the importance of the proposed APA loss during the update of the synthetic images. Fig. 4 demonstrates that the optimal performance is achieved when $\alpha_{1}=1e5$. Incrementally increasing $\alpha_{1}$ improves performance up to 62.29% at $\alpha_{1}=1e5$. However, further increases in $\alpha_{1}$ subsequently degrade performance.
Figure 4: Ablation study of $\alpha_{1}$.
Figure 5: Ablation study of $K_{MSR}$.
The $K_{APA}$ in APA is the upper limit on the number of the Gaussian distributions used for priors generation. Tab. 5 displays the ablation study for different values of $K_{APA}$. The best accuracy is achieved when $K_{APA}=5$.
Priors Distribution Top-1
GMM 62.29
Laplace 62.16
Real 63.19
Table 4: Effect of priors type.
$K_{APA}$ Top-1
1 61.13
3 61.52
5 62.29
7 61.53
9 61.05
Table 5: Effect of $K_{APA}$.
S Top-1
00 61.96
$L/2$ 62.29
Table 6: Effect of $S$.
w. $\frac{l}{L}$ Top-1
62.29
$\times$ 61.32
Table 7: Effect of $\frac{l}{L}$.
The $K_{MSR}$ in MSR is the upper limit on the number of patches. Fig. 5 demonstrates that when $K_{MSR}=4$, the optimal performance is achieved. Using $K_{MSR}$ larger than 4 will hurt the accuracy. We consider this due to limited patch resolution if using a too large $K_{MSR}$.
Analysis of APA loss. Here, we conduct the ablation study by considering the $S$ and scale $\frac{l}{L}$ in Eq. 7. Tab. 7 presents the effect of varying $S$ in Eq. 7. If applying APA loss to all blocks ($S=0$), the top-1 accuracy decreases to 61.96%. From Tab. 7, it can be seen that absorbing the scale $\frac{l}{L}$ in Eq. 7 presents 0.97% performance gains.
5 Limitations
We further discuss some limitations of the proposed SPDFQ, which will guide future research directions. First, although SPDFQ shows substantial performance improvement, a performance gap between SPDFQ and real data remains challenging. Further enhancing semantic alignment between synthetic and real images could help close this gap, highlighting the need for a more precise semantics prompting method. Second, SPDFQ currently lacks a theoretical foundation. Theoretical exploration would be beneficial for gaining more insight into the DFQ for ViTs.
6 Conclusion
In this paper, we investigate DFQ specifically designed for ViTs. We first identify that synthetic images generated by existing methods suffer from the deficient semantics issue, where the semantics deviate from the real images, leading to limited performance. To address this, we propose a novel DFQ method, dubbed SPDFQ, which introduces three key innovations: Attention Priors Alignment (APA), Multi-Semantic Reinforcement (MSR), and Softlabel Learning (SL). APA encourages high responses in attention maps to prompt semantics in synthetic images. MSR promotes efficient parameterization and semantic diversity. SL enables the generation of complex semantics in synthetic images and accommodates images augmented by MSR. Extensive experiments across various ViT models demonstrate the effectiveness of SPDFQ.

### v5

@@
-Semantics Prompting Data-Free Quantization for Low-Bit Vision Transformers
+Semantic Alignment and Reinforcement for Data-Free Quantization of Vision Transformers
 Abstract
-Data-free quantization (DFQ), which facilitates model quantization without real data to address increasing concerns about data security, has garnered significant attention within the model compression community.
-Recently, the unique architecture of vision transformers (ViTs) has driven the development of specialized DFQ techniques.
-However, we observe that the synthetic images from existing methods suffer from the deficient semantics issue compared to real images, thereby compromising performance.
-Motivated by this, we propose SPDFQ, a Semantics Prompting Data-Free Quantization method for ViTs.
-First, SPDFQ incorporates Attention Priors Alignment (APA), which uses randomly generated attention priors to enhance the semantics of synthetic images.
-Second, SPDFQ introduces Multi-Semantic Reinforcement (MSR), which utilizes localized patch optimization to prompt efficient parameterization and diverse semantics in synthetic images.
-Finally, SPDFQ employs Softlabel Learning (SL), where soft learning targets are adapted to encourage more complex semantics and accommodate images augmented by MSR.
-Experimental results demonstrate that SPDFQ significantly outperforms existing methods.
-For instance, SPDFQ achieves a 15.52% increase in top-1 accuracy on ImageNet for W4A4 ViT-B [^1] The code will be made publicly available upon acceptance..
-Figure 1: Feature visualization using $t$-SNE [63].
-Each shape (circle, star, triangle) represents a category.
-The red dashed circles highlight the features of our APA and real images.
-The visualized features are activations extracted before the classification head of DeiT-S.
-The features from PSAQ-ViT [43] deviate significantly from those of the real images, suggesting deficient semantics.
-While the features from our APA are more closely aligned with those of real images, suggesting improved semantics.
+Data-free quantization (DFQ) enables model quantization without accessing real data, addressing concerns regarding data security and privacy.
+With the growing adoption of Vision Transformers (ViTs), DFQ for ViTs has garnered significant attention.
+However, existing DFQ methods exhibit two limitations: (1) semantic distortion, where the semantics of synthetic images deviate substantially from those of real images, and (2) semantic inadequacy, where synthetic images contain extensive regions with limited content and oversimplified textures, leading to suboptimal quantization performance.
+To address these limitations, we propose SARDFQ, a novel Semantics Alignment and Reinforcement Data-Free Quantization method for ViTs.
+To address semantic distortion, SARDFQ incorporates Attention Priors Alignment (APA), which optimizes synthetic images to follow randomly generated structure attention priors.
+To mitigate semantic inadequacy, SARDFQ introduces Multi-Semantic Reinforcement (MSR), leveraging localized patch optimization to enhance semantic richness across synthetic images.
+Furthermore, SARDFQ employs Soft-Label Learning (SL), wherein multiple semantic targets are adapted to facilitate the learning of multi-semantic images augmented by MSR.
+Extensive experiments demonstrate the effectiveness of SARDFQ, significantly surpassing existing methods.
+For example, SARDFQ improves top-1 accuracy on ImageNet by 15.52% for W4A4 ViT-B [^1] The code is at https://github.com/zysxmu/SARDFQ..
+(a) The t-SNE[68] visualization of the penultimate-layer features (extracted by DeiT-S) of synthetic images.
+Each marker (circle, star, triangle) represents a distinct category.
+The red dashed circles highlight the features extracted from our APA and real images.
+Notably, the features produced by PSAQ-ViT[47] exhibit substantial deviation from those of real images, indicating semantic distortion.
+In contrast, our APA yields features that more closely align with those of real images, suggesting improved semantics alignment.
+(b) The images of PSAQ-ViT and PSAQ-ViT V2 exhibit numerous dull regions with limited content and simplified textures, reflecting semantic inadequacy.
+In comparison, our SARDFQ generates images with greater diversity in both content and texture, demonstrating enhanced semantics.
+Figure 1: Illustration of (a) semantic distortion and (b) semantic inadequacy.
 1 Introduction
-Vision transformers (ViTs) [19] have attracted widespread interest from both academia and industry [33, 26] due to their superior performance across various vision tasks [62, 7, 1, 52].
-However, the substantial computational demands and high memory requirements of ViTs pose significant challenges for deployment in resource-limited environments [61, 27, 29, 75, 42, 51, 11].
-To address this challenge, model quantization [36] has emerged as a promising solution, reducing model complexity by enabling a low-bit representation of weights and activations.
-Nevertheless, most existing quantization methods rely on access to the original training data, raising concerns regarding data privacy and security.
-This data dependency restricts their application in data-sensitive environments [6, 77, 24, 70].
-To address this limitation, data-free quantization (DFQ) has been developed to quantize models without accessing real data, attracting considerable attention [69, 74, 14].
-Most existing DFQ methods are tailored to convolutional neural networks (CNNs) and are not directly applicable to ViTs.
-These methods typically utilize batch normalization statistics (BNS), which capture the distribution of real data, to synthesize in-distribution synthetic data [6, 69, 74, 77].
-However, BNS is unavailable for ViTs that use layer normalization (LN) to dynamically compute distribution statistics during inference [43].
-Recently, several works have been proposed to accommodate the unique structure of ViTs [43, 44, 58, 16, 30].
+Vision Transformers (ViTs) [20, 34, 27] have demonstrated remarkable success across various computer vision tasks [67, 7, 1, 56].
+However, their high computational cost and substantial memory footprint hinder deployment in resource-constrained environments [66, 28, 30, 81, 46, 55, 11].
+To address these limitations, quantization [39] has emerged as a promising solution, which reduces model complexity by converting full-precision weight and activations into low-bit representations.
+Traditional quantization methods typically require access to the original training dataset, which raises data privacy and security concerns [6, 83, 25, 76].
+As a result, data-free quantization (DFQ) has gained increasing attention, allowing quantization without the need for real data [75, 80, 14].
+However, most existing DFQ methods are designed specifically for convolutional neural networks (CNNs) and are not directly applicable to vision transformers (ViTs).
+These methods generally rely on batch normalization statistics (BNS), which capture the distribution of real data, to synthesize in-distribution synthetic data [6, 75, 80, 83].
+Yet, BNS is unavailable for ViTs, which use layer normalization (LN) to dynamically compute distribution statistics during inference [47].
+Recently, several DFQ methods have been proposed for ViTs [47, 48, 62, 16, 31].
 For example, PSAQ-ViT [47] introduces patch similarity entropy (PSE) loss to optimize Gaussian noise towards usable synthetic images.
-Nevertheless, we observe that the current synthetic images suffer from the deficient semantics issue.
-In Fig.
-1, we visualize the penultimate features of the synthetic images using t-SNE [63].
-It can be observed that the features of synthetic images generated by PSAQ-ViT diverge significantly from those of real images, suggesting that these synthetic images are semantically deficient.
-This observation is further supported by the quantitative results in Tab.
-1, where the cosine similarity between synthetic images from PSAQ-ViT and real images is notably low.
-For high-bit quantization, where model capacity is largely retained [39], the performance degradation remains relatively minor even with deficient semantics images [43, 58].
-However, in low-bit quantization, where model capacity is severely damaged and informative images are essential for recovering performance [77, 13], fine-tuning on these deficient semantics images leads to poor generalization to real datasets, resulting in limited performance gains.
-Motivated by this finding, we propose a Semantics Prompting Data-Free Quantization method for ViTs, termed SPDFQ.
+Nevertheless, we observe that existing methods suffer from semantic distortion and inadequacy issues.
+As shown in Fig.
+1(a) [^2] We do not plot the results of images from PSAQ-ViT V2 as they lack specific categories., features of synthetic images generated by PSAQ-ViT deviate significantly from those of real images.
+Tab.
+1 shows that the cosine similarity between synthetic images from PSAQ-ViT and real images is notably low, also indicating significant distortion.
+These results highlight the issue of semantic distortion.
+Moreover, as shown in Fig.
+1(b) [^3] More results are presented in the appendix., synthetic images generated by PSAQ-ViT and PSAQ-ViT V2 exhibit many regions with limited content diversity and overly simplified textures.
+These low-quality dull regions are useless or even detrimental to model learning [31], highlighting the issue of semantic inadequacy.
+Consequently, quantized models trained on such low-quality images suffer from degraded performance.
+Motivated by the above analysis, we propose a novel Semantics Alignment and Reinforcement Data-Free Quantization method for ViTs, termed SARDFQ.
 The overall framework is depicted in Fig.
 2.
-First, SPDFQ introduces Attention Priors Alignment (APA), which employs randomly generated priors to encourage high responses in attention maps, thereby explicitly promoting semantics in synthetic images.
-These priors are generated using the Gaussian Mixture Model (GMM) to simulate diverse attention patterns.
-APA effectively prompts semantics in synthetic images, as validated by both visualization and quantitative analyses.
+To address the semantic distortion issue, SARDFQ introduces Attention Priors Alignment (APA), where synthetic images are optimized to follow structured attention priors generated using Gaussian Mixture Models (GMMs).
+APA effectively aligns the semantics of synthetic images with real images, as validated by both visual and quantitative analyses.
 As shown in Fig.
-1, features of APA exhibit a closer alignment to those of real images, indicating high semantics.
-The quantitative results in Tab.
-1 further confirm that the semantics of APA are more similar to the real images.
-Additionally, SPDFQ incorporates Multi-Semantic Reinforcement (MSR) and Softlabel Learning (SL).
-MSR promotes efficient parameterization and diverse semantics in synthetic images by leveraging localized patch-wise optimization, while SL prompts more complex semantics in synthetic images and accommodates images augmented by MSR through the use of soft learning targets.
-Experimental results across various ViT models demonstrate that SPDFQ presents substantial performance improvements.
-For example, SPDFQ achieves a 15.52% increase in top-1 accuracy on the ImageNet dataset for the W4A4 ViT-B model.
+1(a), features of APA exhibit a closer alignment to those of real images, while quantitative results in Tab.
+1 also confirm that the semantics of APA are more consistent to the real images, indicating enhanced semantics alignment.
+To address the semantic inadequacy issue, SARDFQ incorporates Multi-Semantic Reinforcement (MSR) and Softlabel Learning (SL).
+MSR utilizes localized patch optimization, which encourages different sub-patches of synthetic images to capture various semantics, reinforcing the rich semantics across images.
+SL applies multiple semantic targets to accommodate the learning of multi-semantic images augmented by MSR.
+As shown in Fig.
+1(b), synthetic images after applying MSR exhibit greater diversity in content and texture, providing reinforced semantics.
+Experimental results across various ViT models and tasks demonstrate that SARDFQ presents substantial performance improvements.
+For example, SARDFQ achieves a 15.52% increase in top-1 accuracy on the ImageNet dataset for the W4A4 ViT-B model.
 2 Related Works
@@
 However, the model complexity is still unsatisfactory [47].
+Figure 2: SARDFQ Framework overview: Attention Priors Alignment (APA) employs randomly generated attention priors to improve semantics alignment.
+Multi-Semantic Reinforcement (MSR) learns the different regions of synthetic images with various semantics to enhance overall semantic richness.
+Meanwhile, Softlabel Learning (SL) adopts multiple semantic targets to ensure consistent learning of multi-semantic images augmented by MSR.
 2.2 Network Quantization
@@
 Data-driven quantization can be roughly divided into two categories: quantization-aware training (QAT) and post-training quantization (PTQ).
-QAT is compute-heavy since it re-trains the quantized model with the full training data to retain performance [40, 42, 38, 22, 48, 20, 67].
+QAT is compute-heavy since it re-trains the quantized model with the full training data to retain performance [43, 46, 41, 23, 52, 21, 73, 65].
 PTQ perform quantization with a tiny dataset and a reduced time overhead, harvesting widespread attention [3, 42].
@@
 Data-Free Quantization.
-DFQ quantizes models without accessing real data [24, 14, 41, 2, 13, 57, 37, 32, 15].
+DFQ quantizes models without accessing real data [25, 14, 44, 2, 13, 61, 40, 33, 15, 45].
 Most previous DFQ methods focus on CNN, where the BNS can be adopted as the regularization term [6, 80].
@@
 Thus, patch similarity entropy (PSE) loss is proposed to optimize the Gaussian noise towards real-like images by making them showcase heterogeneous patches.
-Based on PSAQ-ViT, PSAQ-ViT V2 [44] further introduces a adversarial learning strategy [23].
+Based on PSAQ-ViT, PSAQ-ViT V2 [48] further introduces an adversarial learning strategy [24].
 [62] incorporates contrastive learning and proposes an iterative generation-quantization PTQ-based DFQ method.
 [31] proposes a sparse generation method to remove noisy and hallucination backgrounds in synthetic images.
-Figure 2: Framework of the proposed SPDFQ.
-Attention Priors Alignment (APA) employs randomly generated attention priors to explicitly promote semantics in synthetic images.
-Multi-Semantic Reinforcement (MSR) prompts the different parts of synthetic images with various semantics.
-Softlabel Learning (SL) leverages soft learning to prompt more complex semantics in synthetic images.
+SynQ [37] introduces an overall framework for DFQ by integrating a low-pass filter, class activation map alignment, and difficult samples learning.
+For a more comprehensive DFQ survey, we invite the reader to see [36].
 3 Method
@@
 $\mathbf{x}_{q}$ $=\text{clip}\left(\left\lfloor\frac{\mathbf{x}}{\Delta}\right\rceil+z,0,2^{b}-1\right),\bar{\mathbf{x}}=\Delta\cdot\left(\mathbf{x}_{q}-z\right),$ (1)
-[…]
+where $\left\lfloor\cdot\right\rceil$ denotes rounding to the nearest integer, and clip limits the value to $0,2^{b}-1]$.
+Here, $\Delta$ and $z$ are the scale factor and zero-point, respectively.
+For the log2 quantizer:
 $\mathbf{x}_{q}$ $=\text{clip}\left(\left\lfloor-\log_{2}\frac{\mathbf{x}}{\Delta}\right\rceil,0,2^{b}-1\right),\bar{\mathbf{x}}=\Delta\cdot 2^{-\mathbf{x}_{q}}.$ (2)
@@
 DFQ methods parameterize synthetic images and optimize them toward real-like images with a pre-trained full-precision model $F$.
-Given a image $\tilde{\bm{I}}$ initialized from Gaussian noise, the one-hot loss [68] is introduced to prompt label-related semantics:
+Given a image $\tilde{\bm{I}}$ initialized from Gaussian noise, the one-hot loss [74] is introduced to learm label-related semantics:
 $\mathcal{L}^{\text{OH}}(\tilde{\bm{I}})=CE(F(\tilde{\bm{I}}),c)$ (3)
@@
 Moreover, total variance (TV) [76] loss is a smoothing regularization term to improve the image quality:
+$\mathcal{L}^{\text{TV}}(\tilde{\bm{I}})=\iint|\nabla\tilde{\bm{I}}(\tau_{1},\tau_{2})|d\tau_{1}d\tau_{2}.$ (4)
+where $\nabla\tilde{\bm{I}}(\tau_{1},\tau_{2})$ denotes the gradient at $\tilde{\bm{I}}$ at $(\tau_{1},\tau_{2})$.
 […]
-where $\nabla\tilde{\bm{I}}(\tau_{1},\tau_{2})$ denotes the gradient at $\tilde{\bm{I}}$ at $(\tau_{1},\tau_{2})$.
-To perform DFQ for ViTs, PSAQ-ViT [43] proposes patch similarity entropy (PSE) loss.
-It first compute patch similarity $\bm{\Gamma}_{l}[i,j]=\frac{u_{i}\cdot u_{j}}{||u_{i}||||u_{j}||}$, where $u_{i},u_{j}$ are feature vectors of MHSA outputs in $l$-th block and |||||$\cdot$||| | ⋅ | | denotes the $l_{2}$ norm.
-Then, it estimates the density function $\hat{f_{l}}(x)=\frac{1}{Mh}\sum_{m=1}^{M}K\left(\frac{x-x_{m}}{h}\right)$, where $K(\cdot)$ is a normal kernel, $h$ is the bandwidth, and $x_{m}$ is the kernel center derived from $\bm{\Gamma}_{l}$.
-Finally, the PSE loss is defined as:
-[…]
+$\mathcal{L}^{\text{PSE}}(\tilde{\bm{I}})=\sum_{l=1}^{L}\int\hat{f_{l}}(x)\log\left[\hat{f_{l}}(x)\right]dx,$ (5)
 where $L$ is the block number of the model.
 3.2 Observations
-Although existing DFQ methods of ViTs have made some progress, a significant performance gap remains between synthetic and real images.
-For example, as shown in Tab.
-2, the W4A4 ViT-B fine-tuned on real images yields 68.16%, whereas PSAQ-ViT only achieves 36.32%.
-To dive into a deeper analysis, we visualize the features of the synthetic images before the classification head using t-SNE [63] in Fig.
-1.
-Generally, the penultimate feature represents the semantics of the input [72, 4, 54, 71].
-It is evident that features of PSAQ-ViT diverge significantly from those of real samples, indicating that the synthetic images fail to capture the real semantics distribution.
-We refer to this phenomenon as the deficient semantics issue.
-In Tab.
-1, we quantitatively measure the semantics using the average cosine similarity (ranging from $-1$ to $1$).
-We also report the intra-class similarity within real images as an approximate upper bound for comparison.
-These results further highlight the deficient semantics of PSAQ-ViT.
-For example, for Class 1, the intra-class similarity within real images is 0.68, whereas PSAQ-ViT only presents 0.44, suggesting a low similarity between synthetic and real images.
-Similarly, for Class 3, while the real images exhibit a similarity of 0.41, PSAQ-ViT achieves only 0.31.
-Both visualization and quantitative results reveal a pronounced deficient semantics.
-Consequently, quantized models fine-tuned on these semantically deficient synthetic images cannot well generalize to the real dataset, resulting in limited performance.
 Classes Method Real PSAQ-ViT [47] APA (Ours)
@@
 For real images, the similarity is measured within the class itself, while for PSAQ-ViT and APA, the similarity is measured between synthetic and real images of the same class.
-The results show that APA achieves higher similarity than PSE loss, indicating improved semantics.
-3.3 Solutions
-In the following, we introduce the proposed SPDFQ to prompt semantics in synthetic images.
-As illustrated in Fig.
-2, SPDFQ introduces three contributions: Attention Priors Alignment (APA), Multi-Semantic Reinforcement (MSR), and Softlabel Learning (SL).
+The results show that APA achieves higher similarity than PSE loss, indicating aligned semantics.
+Existing DFQ methods have made significant progress.
+However, through carefully analyzing the synthetic images, we reveal that these images suffer from issues of semantic distortion and semantic inadequacy, both of which hinder further advancement in the DFQ of ViTs.
+The semantic distortion issue refers to the significant divergence between the semantics of synthetic images and real images.
+To demonstrate this, we visualize the features of synthetic and real images in Fig.
+1(a).
+Note that the penultimate feature typically is regarded to represent the semantics of the input [78, 4, 58, 77], It is clear that the features of PSAQ-ViT diverge significantly from real images, suggesting that these images fail to capture the true semantic distribution of real data.
+Tab.
+1 quantitatively measures the semantics using the average cosine similarity (ranging from 1-1 to $1$).
+We also report the intra-class similarity within real images as an approximate upper bound for comparison.
+This result further supports the observation of low similarity between synthetic images from PSAQ-ViT and real images.
+For instance, for Class 1, the intra-class similarity within real images is 0.68, while PSAQ-ViT achieves only 0.44.
+The semantic inadequacy issue refers to the presence of dull regions in synthetic images, which contain redundant or non-semantic content [31, 35], hindering the model’s learning process.
+As indicated in [14, 13], a diverse content and textures generally suggests rich information.
+As shown in Fig.
+1(b), many regions of synthetic images generated by PSAQ-ViT and PSAQ-ViT V2 exhibit a lack of diversity in content, with overly simplified textures.
+Specifically, the central region of PSAQ-ViT images only contains faint object structures, while PSAQ-ViT V2 images appear excessively smoothed and indistinct.
+For high-bit quantization, where model capacity is largely retained [42], the performance degradation remains relatively minor even using semantic distorted and inadequate images [47, 62].
+However, in low-bit quantization, where model capacity is severely damaged and informative images are essential for recovering performance [83, 13], fine-tuning on these poor-quality images leads to inferior generalization to real datasets, resulting in limited performance.
+For example, as shown in Tab.
+2, the W4A4 ViT-B fine-tuned on real images yields 68.16%, whereas PSAQ-ViT only achieves 36.32%.
+Figure 3: Comparison between attention maps.
+3.3 Semantics Alignment and Reinforcement Data-Free Quantization
+In the following, we introduce the proposed SARDFQ, whose framework is illustrated in Fig. 2.
 3.3.1 Attention Priors Alignment
-The first step of SFDFQ in prompting semantics leverages the self-attention mechanism of ViTs.
-Specifically, the self-attention mechanism extracts and integrates semantic information by allocating more attention to tokens rich in semantics and less to tokens with minimal semantics [19, 28, 10].
-Thus, if a synthetic image is semantically rich, its attention maps should exhibit high responses.
-Based on this characteristic, we propose Attention Priors Alignment (APA), which is designed to promote the semantics of synthetic images by explicitly encouraging high responses in images’ attention maps.
-For a synthetic image $\tilde{\bm{I}}$, we first extract its attention maps of the $h$-th head in the $l$-th block, denoted by $\mathbf{A}_{l,h}\in\mathbb{R}^{N\times N}$, by inputting the image into a pre-trained full-precision model, where $N$ represents the total number of tokens.
+In ViTs, the self-attention mechanism encodes semantic correlations between image regions, where high-response areas in attention maps strongly correlate with semantic-discriminative content [20, 29, 10].
+However, existing DFQ methods overlook this intrinsic property in the generation process.
+As a result, as shown in Fig.
+3, synthetic images often exhibit disordered and unnatural attention patterns, with attention maps either overly diffuse or misaligned toward peripheral regions.
+This undermines their ability to preserve semantic-discriminative content, causing semantic distortion, as demonstrated in Fig.
+1(a) and Tab.
+1.
+In response, we propose Attention Priors Alignment (APA), which improves semantics alignment by optimizing synthetic images to follow randomly generated structure attention priors.
+Specifically, given a synthetic image $\tilde{\bm{I}}$, we first obtain its attention maps in the $h$-th head of the $l$-th block, denoting as $\mathbf{A}_{l,h}\in\mathbb{R}^{N\times N}$, where $N$ represents the total number of tokens.
 In DeiT, the attention of the classification token toward other tokens serves as the indicator for semantic versus non-semantic parts [5].
-Accordingly, we extract $\mathbf{A}^{c}_{l,h}\in\mathbb{R}^{1\times(N-1)}$ from $\mathbf{A}_{l,h}$, representing the attention of the classification token to all tokens except itself.
+Thus, we extract $\mathbf{A}^{c}_{l,h}\in\mathbb{R}^{1\times(N-1)}$ from $\mathbf{A}_{l,h}$, representing the attention of the classification token to all tokens except itself.
 We then randomly generate attention priors $\tilde{\mathbf{A}}_{l,h}$, whose generation is detailed in the next part, and align $\mathbf{A}^{c}_{l,h}$ with $\tilde{\mathbf{A}}_{l,h}$ by:
@@
 6 across these deeper blocks:
-$\mathcal{L}^{\text{APA}}(\tilde{\bm{I}})=\sum_{l=S}^{L}\sum_{h=1}^{H}\frac{l}{L}\mathcal{L}_{l,h}(\tilde{\bm{I}}),$ (7)
+[…]
 where $S$ is a pre-given hyper-parameter denoting the start of deep blocks and are experimentally set $S=\frac{L}{2}$.
-The proposed APA loss encourages synthetic images to contain semantically meaningful regions with high responses in attention maps, thereby prompting the semantics of the generated images.
 Figure 4: Examples of generated attention priors.
 Attention Priors Generation.
-To generate attention priors $\tilde{\mathbf{A}}_{l,h}$, we employ GMM for it is the most commonly used distribution with high flexibility.
+To generate attention priors $\tilde{\mathbf{A}}_{l,h}$, Gaussian Mixture Models (GMMs) are employed as it is the most commonly used distribution with high flexibility.
 ViTs utilize different attention heads to capture diverse patterns and learn varied, informative representations [8].
 Thus, we use distinct GMMs for each head.
-Note that the goal here is not to replicate real attention maps precisely, but to generate simulated attention priors.
-These priors are used to explicitly encourage semantically meaningful regions within synthetic images, prompting high responses in attention maps to enhance semantics.
+Note that the goal here is not to replicate real attention maps precisely, but to generate simulated structure attention priors to guide the learning of synthetic images.
 In particular, we first initialize an all zero matrix $\tilde{\mathbf{P}}\in\mathbb{R}^{H\times W}$, where $H=W=\sqrt{N-1}$ for DeiT, $H=W=\sqrt{N}$ for Swin.
@@
 Then, we generate $k$ two-dimensional Gaussian distributions, where $k$ is randomly sampled from $1\sim K_{APA}$ and $K_{APA}$ is set to $5$ in all experiments.
-Each Gaussian has its mean and covariance [^2] The pseudo code is detailed in supplementary materials..
+Each Gaussian has its mean and covariance [^4] The pseudo code is detailed in appendix..
 Consequently, the matrix element at the $i$-th row and $j$-th column, $\tilde{\mathbf{P}}[i,j]$, is determined by:
@@
 Then, $\tilde{\mathbf{P}}$ is normalized by:
-$\tilde{\mathbf{P}}_{n}=\frac{\tilde{\mathbf{P}}}{\sum\tilde{\mathbf{P}}}*(1-x).$ (9)
+$\tilde{\mathbf{P}}_{n}=\frac{\tilde{\mathbf{P}}}{\sum\tilde{\mathbf{P}}}\cdot(1-x).$ (9)
 Here, for DeiT that incorporates the classification token, $x$ is randomly sampled from a uniform distribution $U(0,1)$, representing the proportion of the attention score that the classification token allocates to itself.
@@
 $\tilde{\mathbf{A}}_{l,h}=\text{flatten}(\tilde{\mathbf{P}}_{n}).$ (10)
-Fig.
-3 displays examples of generated attention priors.
-Although simple, APA effectively enhances semantics, as validated by both visualization and quantitative results.
+Fig. 4 displays examples of generated attention priors.
+Discussion.
+APA prevents attention disorder, ensuring that synthetic images exhibit more coherent and natural attention patterns, as demonstrated in Fig.
+3.
+As a result, APA selectively enhances the responses in certain regions, effectively emphasizing semantic-discriminative regions within synthetic images and thus prompting the discriminative features.
+Although simple, APA enables synthetic images to align better with the real semantics, as validated by both visual and quantitative evaluations.
 As shown in Fig.
-1, compared to PSAQ-ViT, features after applying the APA loss are closer to the real images, indicating improved semantics.
+1(a), compared to PSAQ-ViT, features obtained after applying the APA loss are more closely aligned with real images, indicating better semantic alignment.
 The quantitative results in Tab.
-1 further support that APA achieves higher semantics.
-For instance, in Class 1, the intra-class similarity of PSAQ-ViT is 0.44, whereas APA achieves an improved similarity of 0.64.
+1 further support that APA achieves superior semantic alignment.
+For example, in Class 1, the intra-class similarity for PSAQ-ViT is 0.44, whereas APA achieves a higher similarity of 0.64.
 3.3.2 Multi-Semantic Reinforcement
-Current methods [43, 44, 58] optimize the entire synthetic image, restricting it to capture only the semantics of a single object type.
-However, this has two limitations.
-First, real images naturally exhibit low-rank properties, such as spatial similarity between nearby pixels [31].
-Optimizing the entire images overlooks this structural regularity, easily leading to synthetic images with noisy or spurious regions that convey meaningless or misleading semantics [30].
-This results in inefficient parameterization, where semantic contents available for model training are limited [34].
-Second, real images often contain multiple objects with varied semantics.
-Constraining synthetic images to a single object type limits their semantic richness and diversity.
-Motivated by the above analysis, we propose Multi-Semantic Reinforcement (MSR), which leverages localized patch optimization to promote efficient parameterization and diverse semantics.
+Current DFQ methods for ViTs [47, 48, 62] optimize images through global optimization, treating the entire image as a single semantic unit.
+However, this is affected by low-rank structural regularity [32], where adjacent pixels exhibit strong similarity, leading to dull regions with redundant or non-semantic content for model learning [35].
+This issue is further exacerbated by the tokenization mechanism in ViTs, as processing images in fixed-size patches make dull regions increase at the patch level [19, 69].
+Consequently, as shown in Fig.
+1(b), synthetic images generated by existing methods exhibit large dull regions, resulting in semantic inadequacy [35].
+In response, we propose Multi-Semantic Reinforcement (MSR), which applies localized patch optimization to enhance semantic richness by learning local patches with distinct semantics.
 Specifically, for a synthetic image $\tilde{\bm{I}}$, instead of feeding only the entire image, we also feed its patches and optimize them individually.
@@
 These $m$ patches are then cropped and resized to match the model’s input dimensions:
-$\{\tilde{\bm{I}}^{i}_{MSR}\}_{i=1,\cdots,m}=\text{resize}\big{(}\text{crop}_{m}(\tilde{\bm{I}})\big{)},$ (11)
+$),\{\tilde{\bm{I}}^{i}_{MSR}\}_{i=1,\cdots,m}=\text{resize}\big(\text{crop}_{m}(\tilde{\bm{I}}$$\big)$, (11)
 where $\text{crop}_{m}(\cdot)$ crops $m$ non-overlapping patches from input, and $\text{resize}(\cdot)$ is the resize function.
 Each patch, denoted as $\tilde{\bm{I}}^{i}_{MSR}$, is treated as a new image with an assigned semantic target $c^{i}$.
-For simplicity, the loss function used to learn label-related semantics in both $\tilde{\bm{I}}^{i}_{MSR}$ and $\tilde{\bm{I}}$ is presented in Sec.
-3.3.3.
 Note that the gradient is backpropagated only to update the corresponding patch in the original image, leaving the rest of the image unaffected.
-By leveraging localized patch optimization, MSR effectively promotes efficient parameterization and diverse semantics in synthetic images.
-First, MSR reduces noisy and non-semantic regions and provides more distinct semantic samples, i.e., $\{\tilde{\bm{I}}^{i}_{MSR}\}_{i=1,\cdots,m}$, for model training, achieving efficient parameterization.
-Second, by assigning different semantic targets to individual cropped patches, MSR transforms synthetic images into composites of multiple semantic objects rather than a single uniform entity, thereby enhancing overall semantic diversity.
-Unlike traditional cropping used in data augmentation, which aims to improve classification robustness by training with cropped patches labeled with the original class, MSR aims to prompt efficient parameterization and semantic diversity within synthetic images, ultimately enabling accurate data-free quantization.
-3.3.3 Softlabel Learning
+Softlabel Learning.
 The one-hot loss (Eq.
-3) optimizes synthetic images until their confidence for the target class approaches 1.
-However, this leads to overfitting on a specific class, limiting semantic richness [77].
-Moreover, Eq.
-3 only learns the semantics of the target class.
-making it unsuitable for $\tilde{\bm{I}}$ under MSR, as its patches $\{\tilde{\bm{I}}^{i}_{MSR}\}_{i=1,\cdots,m}$ contain distinct semantics.
-To address this, we propose Softlabel Learning (SL), which softens the learning objective to promote more complex semantics and accommodate images augmented by MSR.
+3) only learns the semantics of the target class, making it unsuitable for $\tilde{\bm{I}}$ under MSR, as its patches $\{\tilde{\bm{I}}^{i}_{MSR}\}_{i=1,\cdots,m}$ contain distinct semantics.
+In response, we propose Softlabel Learning (SL), which applies multiple semantic targets to accommodate the learning of images augmented by MSR.
 Specifically, we first sample $Z\in\mathbb{R}^{C}\sim U(0,1)$, then modify its values by:
-$\begin{cases}Z[c^{i}]\sim U(\epsilon_{1},\epsilon_{2}),&\text{for }\tilde{\bm{I}}^{i}_{MSR},\\ Z[c^{1},\ldots,c^{m}]\sim U(\epsilon_{1},\epsilon_{2}),&\text{for }\tilde{\bm{I}},\end{cases}$
-where $U(\epsilon_{1},\epsilon_{2})$ denotes the uniform distribution over the interval $[\epsilon_{1},\epsilon_{2}]$, $m$ is the number of patches determined in MSR, and $\epsilon_{1}$ and $\epsilon_{2}$ control the softness, both empirically set consistently to 5 and 10 in all experiments.
+[…]
+where $U(\epsilon_{1},\epsilon_{2})$ denotes the uniform distribution over the interval $\epsilon_{1},\epsilon_{2}]$, $m$ is the number of patches determined in MSR, and $\epsilon_{1}$ and $\epsilon_{2}$ control the softness, both empirically set consistently to 5 and 10 in all experiments.
 The soft target is defined as $T_{s}=\text{softmax}(Z)$, and SL loss is:
@@
 where $SCE(\cdot,\cdot)$ is soft cross entropy and $F(\cdot)$ returns the predicted probability for its input.
-For $\tilde{\bm{I}}^{i}_{MSR}$, SL aligns its predicted results with a “soft label” rather than a traditional one-hot label, mitigating overfitting and promoting the generation of semantically complex content.
-For $\tilde{\bm{I}}$, SL provides a suitable target for learning images with multiple semantic targets.
-Unlike the Soft Inception loss [77], which focuses only on the primary class probability, SL imposes a distribution across all classes, offering a comprehensive learning objective.
-Additionally, SL is compatible with MSR, where images contain multiple semantic targets.
-In contrast, the Soft Inception loss is designed for single-class scenarios and is therefore incompatible.
+SL facilitates smooth learning between across semantic targets, ensuring that MSR-enhanced images receive consistent supervision rather than conflicting supervision.
+Discusion.
+By leveraging localized patch optimization, MSR ensures that each patch contributes unique semantics, forcing synthetic images to capture diverse features rather than being dominated by large homogeneous dull regions.
+As demonstrated by the richer and more diverse content and textures in Fig.
+1(b), MSR effectively reduces dull regions and enhances semantic richness in synthetic images.
+Moreover, MSR transforms synthetic images into composites of multiple semantic objects rather than a single unit, thereby providing more distinct semantic samples, i.e., $\{\tilde{\bm{I}}^{i}_{MSR}\}_{i=1,\cdots,m}$, for model training.
+Unlike traditional cropping used in data augmentation, which aims to improve classification robustness by training with cropped patches labeled with the original class, MSR aims for semantic richness within synthetic images, ultimately enabling accurate data-free quantization.
 3.4 Overall Pipeline
 The overall pipeline consists of two stages: data synthesis and quantized network learning.
-The first stage synthesizes synthetic images.
+The first stage uses the proposed SARDFQ to produce synthetic images.
 The second stage fine-tunes the quantized model using the generated synthetic images.
@@
 In the data synthesis stage, we combine the proposed APA loss of Eq. 7, SL loss of Eq. 12, and TV loss of Eq. 4 to formulate the objective function as follows:
-$\mathcal{L}_{G}(\tilde{\bm{I}})$ $=\alpha_{1}\mathcal{L}^{\text{APA}}(\tilde{\bm{I}})+{\cal L}^{\text{SL}}(\tilde{\bm{I}})+0.05\mathcal{L}^{\text{TV}}(\tilde{\bm{I}}).$ (13)
+[…]
 where $\alpha_{1}$ is hyperparameters and is determined by grid-search. Note that the weight of TV loss is fixed to 0.05, following [47], to avoid a cumbersome hyperparameter search.
@@
 4.1 Implementation Details
-Models and Datasets.
-We evaluate the performance of the proposed SPDFQ across various ViTs on the image classification task, including ViT-S/B [19], DeiT-T/S/B [62], and Swin-S/B [50].
-The dataset is ImageNet [59], which contains 1.28M training images and 50K validation images across 1,000 classes.
+Models and Tasks.
+We evaluate the performance of SARDFQ by test quantized ViT-S/B [20], DeiT-T/S/B [67], and Swin-S/B [54] on the classification task using ImageNet [63].
 The pre-trained models are downloaded from the timm library.
-Model W/A Real Gaussian noise PSAQ-ViT [43] PSAQ-ViT V2 [44] SMI [30] SPDFQ (Ours)
-ViT-S (81.39) 4/4 66.57 6.02 47.24 41.53 24.3329.41 50.32
+In appendix, we further provide results on detection and segmentation tasks.
+Model W/A Real Gaussian noise PSAQ-ViT [47] PSAQ-ViT V2 [48] SMI [31] SARDFQ (Ours)
+ViT-S
+4/4 66.57 6.02 47.24 41.53 24.3329.41 50.32
 5/5 76.69 36.77 71.59 68.41 61.3365.19 74.31
 6/6 79.46 61.20 77.20 74.76 72.9572.46 78.40
-ViT-B (84.54) 4/4 68.16 0.15 36.32 26.32 35.2719.67 51.84
+ViT-B
+4/4 68.16 0.15 36.32 26.32 35.2719.67 51.84
 5/5 79.21 4.16 68.48 67.95 67.5357.13 70.70
 6/6 81.89 55.18 76.65 71.87 76.3369.82 79.16
-DeiT-T (72.21) 4/4 56.60 17.43 47.75 30.20 30.1413.18 52.06
+DeiT-T
+4/4 56.60 17.43 47.75 30.20 30.1413.18 52.06
 5/5 67.09 43.49 64.10 55.16 56.4439.35 66.41
 6/6 69.81 56.23 68.37 62.77 64.0344.39 69.73
-DeiT-S (79.85) 4/4 68.46 20.89 58.28 45.53 42.7711.71 62.29
+DeiT-S
+4/4 68.46 20.89 58.28 45.53 42.7711.71 62.29
 5/5 75.06 41.06 71.90 63.14 62.8829.13 74.06
 6/6 77.87 65.63 75.85 68.85 71.6537.69 77.31
-DeiT-B (81.85) 4/4 77.07 47.20 71.75 66.43 65.3359.04 72.17
+DeiT-B
+4/4 77.07 47.20 71.75 66.43 65.3359.04 72.17
 5/5 79.86 65.46 78.45 76.77 76.7475.33 78.72
 6/6 80.90 62.79 80.00 79.22 78.8177.66 80.15
-Swin-S (83.20) 4/4 78.12 31.92 73.19 65.55 65.85 74.74
+Swin-S
+4/4 78.12 31.92 73.19 65.55 65.85 74.74
 5/5 80.51 52.10 78.15 74.37 75.41 79.56
 6/6 80.60 65.66 79.74 78.50 78.25 80.56
-Swin-B (85.27) 4/4 78.80 30.14 71.84 67.42 65.23 76.42
+Swin-B
+4/4 78.80 30.14 71.84 67.42 65.23 76.42
 5/5 82.51 35.28 78.50 77.20 75.25 80.82
@@
 Comparison methods.
-We compare our SPDFQ against Gaussian noise, real images, and previous methods including SMI [30] and PSAQ-ViT [43] and its subsequent version, PSAQ-ViT V2 [44].
+We compare our SARDFQ against Gaussian noise, real images, and previous methods including SMI [31] and PSAQ-ViT [47] and its subsequent version, PSAQ-ViT V2 [48].
 For a fair comparison, we generate synthetic images using their methods and apply our quantized network learning strategy.
 We use the official code for SMI and PSAQ-ViT to reproduce their images, while PSAQ-ViT V2 is re-implemented by us, as no official code is available.
-[^3] In supplementary materials, we present more comparisons with more methods on W8/A8 and W4/A8 settings.
+[^5] In appendix, we present practical efficiency results and more performance comparisons with more methods such as CLAMP-ViT [62] on W8/A8 and W4/A8 settings.
 Experimental settings.
-All experiments use the PyTorch framework [56] on a single NVIDIA 3090 GPU.
-In the data synthesis stage, synthetic images were initialized with standard Gaussian noise, with 32 synthetic images generated in total.
-The Adam optimizer [35] with $\beta_{1}=0.5$, $\beta_{2}=0.9$ is employed.
-The learning rate is 0.25 for Swin and 0.2 for others, with a total of 1,000 iterations.
-For all models, $K_{APA}$, $K_{MSR}$, $\epsilon_{1}$, and $\epsilon_{2}$ are respectively set to 5 and 4, 5, and 10, with these values are searched using W4A4 DeiT-S.
-The value of $\alpha_{1}$ was determined by grid search for each model.
-Specifically, $\alpha_{1}$ is set to 1e5 for DeiT-T/S, 1e4 for DeiT-B, 100 for ViT-B and Swin-B, 10 for Swin-S, and 1 for ViT-S.
-Although further searching hyperparameters may improve the performance, we use current settings since these already yield superior results.
-In the quantized network learning stage, the Adam optimizer with $\beta_{1}=0.9,\beta_{2}=0.999$ is employed.
-The weight decay is set to 0, and the initialized learning rate is 4e-5, adjusted by the cosine decay strategy.
-The total iteration is set to 100.
-Consistent with [45], quantization parameters were determined by searching for percentile values that minimize quantization error before and after quantization.
-We follow the two-stage optimization strategy in [78] to optimize the quantized network.
-The channel-wise quantizer is adopted for weights and the layer-wise quantizer for activations.
-All matrix multiplications in ViTs are quantized [44, 43, 45].
+All experiments were conducted using the PyTorch framework [60] on a single NVIDIA 3090 GPU.
+In the data synthesis stage, synthetic images were initialized with standard Gaussian noise, generating 32 images in total.
+The Adam optimizer [38] with $\beta_{1}=0.5$, $\beta_{2}=0.9$ was used, with learning rates of 0.25 for Swin and 0.2 for others, and a total of 1,000 iterations.
+For all models, $K_{APA}$, $K_{MSR}$, $\epsilon_{1}$, and $\epsilon_{2}$ were set to 5, 4, 5, and 10, respectively, based on a search with W4A4 DeiT-S.
+The value of $\alpha_{1}$ was determined by grid search for each model: 1e5 for DeiT-T/S, 1e4 for DeiT-B, 100 for ViT-B and Swin-B, 10 for Swin-S, and 1 for ViT-S.
+Although further hyperparameter search may improve performance, the current settings already yield superior results.
+In the quantized network learning stage, following [84], the Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$ was used, with weight decay set to 0 and an initial learning rate of 4e-5, adjusted via cosine decay for 100 iterations.
+A channel-wise quantizer was used for weights, and a layer-wise quantizer for activations, with all matrix multiplications in ViTs quantized [48, 47, 49].
 4.2 Quantization Results
@@
 2.
-The proposed SPDFQ demonstrates consistent improvements across various quantization bit-width configurations, particularly with low bit-width settings.
-Specifically, for ViT-S, SPDFQ improves the performance by 3.08% in the W4/A4 setting, 2.72% in the W5/A5 setting, and 1.20% in the W6/A6 setting.
-For ViT-B, SPDFQ achieves performance gains of 15.52% in the W4/A4 setting, 2.22% in the W5/A5 setting, and 2.51% in the W6/A6 setting.
-Results on DeiT also demonstrate the effectiveness of the proposed SPDFQ.
-For DeiT-T, SPDFQ shows a marked improvement by increasing top-1 accuracy by 4.31% in the W4/A4 setting, 2.31% in the W5/A5 setting, and 1.36% in the W6/A6 setting.
-For DeiT-S, SPDFQ enhances top-1 accuracy by 4.01% in the W4/A4 setting, 2.16% in the W5/A5 setting, and 1.46% in the W6/A6 setting.
-As for DeiT-B, SPDFQ enhances top-1 accuracy by 0.42% in the W4/A4 setting, 0.27% in the W5/A5 setting, and 0.15% in the W6/A6 setting.
-The quantization results of Swin-S/B also affirm the superiority of our SPDFQ in enhancing model accuracy under different quantization configurations.
-In particular, for Swin-S, the proposed SPDFQ increases the accuracy by 1.55% for the W4/A4 setting, 1.41% for the W5/A5 setting, and 0.82% for the W6/A6 setting, respectively.
-When it comes to Swin-B, the proposed SPDFQ increases the accuracy by 4.58% for the W4/A4 setting, 2.32% for the W5/A5 setting, and 1.03% for the W6/A6 setting, respectively.
+Our SARDFQ demonstrates consistent improvements across various quantization bit-width configurations, particularly with low bit-width settings.
+Specifically, for ViT-S, SARDFQ improves the performance by 3.08% in the W4/A4 setting, 2.72% in the W5/A5 setting, and 1.20% in the W6/A6 setting.
+For ViT-B, SARDFQ achieves performance gains of 15.52% in the W4/A4 setting, 2.22% in the W5/A5 setting, and 2.51% in the W6/A6 setting.
+Results on DeiT also demonstrate the effectiveness of SARDFQ.
+For example, on DeiT-T, SARDFQ shows a marked improvement by increasing top-1 accuracy by 4.31% in the W4/A4 setting, 2.31% in the W5/A5 setting, and 1.36% in the W6/A6 setting.
+For DeiT-S, SARDFQ enhances top-1 accuracy by 4.01% in the W4/A4 setting, 2.16% in the W5/A5 setting, and 1.46% in the W6/A6 setting.
+The quantization results of Swin-S/B also affirm the superiority of our SARDFQ in enhancing model accuracy under different quantization configurations.
+In particular, for Swin-S, the proposed SARDFQ increases the accuracy by 1.55% for the W4/A4 setting, 1.41% for the W5/A5 setting, and 0.82% for the W6/A6 setting, respectively.
+When it comes to Swin-B, the proposed SARDFQ increases the accuracy by 4.58% for the W4/A4 setting, 2.32% for the W5/A5 setting, and 1.03% for the W6/A6 setting, respectively.
 4.3 Ablation Study
@@
 Analysis of APA, MSR, and SL.
-We first analyze the effectiveness of the proposed APA (Sec.
-3.3.1), MSR (Sec.
-3.3.2), and SL (Sec.
-3.3.3).
-Experimental results are present in Tab.
-3.
-It can be seen that when the APA and SL are individually added to synthesize fake images, the accuracy increases compared with the baseline.
-Notably, APA significantly boosts the baseline from 51.73% to 60.26%, supporting the analysis in Sec.
-3.3.1 that demonstrates the APA’s effectiveness in enhancing semantics.
-In contrast, applying MSR alone slightly reduces accuracy from 51.73% to 50.75%, indicating that one-hot loss is unsuitable for synthetic images augmented with MSR.
-However, when both MSR and SL are applied, performance rises to 56.08%, suggesting that SL is more compatible with MSR than the one-hot loss.
-Combining APA with either MSR or SL further increases performance.
-For example, applying both APA and SL presents an accuracy of 60.51%.
-When all of the three strategies are applied, the best performance of 62.29% can be obtained.
+We analyze the effectiveness of the proposed APA (Sec.,3.3.1), MSR (Sec.,3.3.2), and SL (Eq.,12) in Tab.,3.
+Adding APA and SL individually to the baseline increases accuracy.
+Notably, APA boosts performance from 51.73% to 60.26%, confirming its effectiveness in aligning semantics (Sec.,3.3.1).
+Applying MSR alone slightly decreases accuracy from 51.73% to 50.75%, indicating that one-hot loss is unsuitable for MSR-augmented synthetic images.
+However, when both MSR and SL are applied, accuracy rises to 56.08%, suggesting SL is more compatible with MSR than one-hot loss.
+Combining APA with either MSR or SL further improves performance.
+For example, APA and SL together yield an accuracy of 60.51%, and when all three strategies are used, the best performance of 62.29% is achieved.
 Analysis of priors distribution.
 Tab.
-5 showcases the results of using other distributions to formulate the attention priors.
+4(a) showcases the results of using other distributions to formulate the attention priors.
 The unevenly distributed GMM and Laplace present comparable performance of 62.29% and 62.16%, respectively.
 Moreover, GMM provides a similar performance to the real’s, indicating it performs well in imitating the patterns of real images.
+Figure 5: Effect of varying (a) $\alpha_{1}$ and (b) $K_{MSR}$.
 Analysis of $\alpha_{1}$, $K_{APA}$, and $K_{MSR}$.
@@
 Fig.
-4 demonstrates that the optimal performance is achieved when $\alpha_{1}=1e5$.
+5(a) demonstrates that the optimal performance is achieved when $\alpha_{1}=1e5$.
 Incrementally increasing $\alpha_{1}$ improves performance up to 62.29% at $\alpha_{1}=1e5$.
 However, further increases in $\alpha_{1}$ subsequently degrade performance.
-Figure 4: Ablation study of $\alpha_{1}$.
-Figure 5: Ablation study of $K_{MSR}$.
 The $K_{APA}$ in APA is the upper limit on the number of the Gaussian distributions used for priors generation.
 Tab.
-5 displays the ablation study for different values of $K_{APA}$.
+4(b) displays the ablation study for different values of $K_{APA}$.
 The best accuracy is achieved when $K_{APA}=5$.
+The $K_{MSR}$ in MSR is the upper limit on the number of patches.
+Fig.
+5(b) demonstrates that when $K_{MSR}=4$, the optimal performance is achieved.
+Using $K_{MSR}$ larger than 4 will hurt the accuracy.
+We consider this due to limited patch resolution if using a too large $K_{MSR}$.
 Priors Distribution Top-1
@@
 Real 63.19
-Table 4: Effect of priors type.
 $K_{APA}$ Top-1
@@
 9 61.05
-Table 5: Effect of $K_{APA}$.
 S Top-1
@@
 $L/2$ 62.29
-Table 6: Effect of $S$.
 w. $\frac{l}{L}$ Top-1
@@
 $\times$ 61.32
-Table 7: Effect of $\frac{l}{L}$.
-The $K_{MSR}$ in MSR is the upper limit on the number of patches.
-Fig.
-5 demonstrates that when $K_{MSR}=4$, the optimal performance is achieved.
-Using $K_{MSR}$ larger than 4 will hurt the accuracy.
-We consider this due to limited patch resolution if using a too large $K_{MSR}$.
+Table 4: Effect of varying (a) priors types; (b) $K_{APA}$; (c) $S$; (d) $\frac{l}{L}$.
 Analysis of APA loss.
@@
 Tab.
-7 presents the effect of varying $S$ in Eq.
+4(c) presents the effect of varying $S$ in Eq.
 7.
@@
 From Tab.
-7, it can be seen that absorbing the scale $\frac{l}{L}$ in Eq.
+4(d), it can be seen that absorbing the scale $\frac{l}{L}$ in Eq.
 7 presents 0.97% performance gains.
 5 Limitations
-We further discuss some limitations of the proposed SPDFQ, which will guide future research directions.
-First, although SPDFQ shows substantial performance improvement, a performance gap between SPDFQ and real data remains challenging.
-Further enhancing semantic alignment between synthetic and real images could help close this gap, highlighting the need for a more precise semantics prompting method.
-Second, SPDFQ currently lacks a theoretical foundation.
-Theoretical exploration would be beneficial for gaining more insight into the DFQ for ViTs.
+We further discuss some limitations of the proposed SARDFQ, which will guide future research directions.
+First, although SARDFQ shows substantial performance improvement, a performance gap between SARDFQ and real data remains challenging, highlighting the need for a stronger semantics alignment and reinforcement method.
+Second, SARDFQ currently lacks a theoretical foundation.
+Future work could establish a theoretical framework for SARDFQ, particularly in understanding how APA and MSR influence synthetic images in a formalized manner.
 6 Conclusion
-In this paper, we investigate DFQ specifically designed for ViTs.
-We first identify that synthetic images generated by existing methods suffer from the deficient semantics issue, where the semantics deviate from the real images, leading to limited performance.
-To address this, we propose a novel DFQ method, dubbed SPDFQ, which introduces three key innovations: Attention Priors Alignment (APA), Multi-Semantic Reinforcement (MSR), and Softlabel Learning (SL).
-APA encourages high responses in attention maps to prompt semantics in synthetic images.
-MSR promotes efficient parameterization and semantic diversity.
-SL enables the generation of complex semantics in synthetic images and accommodates images augmented by MSR.
-Extensive experiments across various ViT models demonstrate the effectiveness of SPDFQ.
+In this paper, we investigate the DFQ method for ViTs.
+We first identify that synthetic images generated by existing methods suffer from semantic distortion and inadequacy issues, and propose SARDFQ to address these issues.
+To mitigate semantic distortion, SARDFQ introduces APA, which guides synthetic images to align with randomly generated structural attention patterns.
+To tackle semantic inadequacy, SARDFQ incorporates MSR.
+MSR optimizes different regions of synthetic images with unique semantics, thereby enhancing overall semantic richness.
+Moreover, SARDFQ employs SL, which adopts multiple semantic targets to ensure seamless learning of images augmented by MSR.
+Extensive experiments on various ViT models and tasks validate the effectiveness of SARDFQ.